OpenStudy (anonymous):

If the following system -15x + 9y = -9 -25x + ky = -12 is consistent, then k is not equal to = ?

HanAkoSolo (jamierox4ev3r):

this user isn't trying to find x and y, if I am understanding the problem correctly. The user is trying to find what k is equal to, assuming that the following system of equations is consistent.

OpenStudy (anonymous):

Yes, i am looking for k.

OpenStudy (anonymous):

OK ! NOW LOOK !: -15x + 9y = -9 -25x + ky = -12 -9/-12 = 3/4 Now do it : 3/4(-25x + ky = -12) We have : -75/4x + 3/4ky = -9 -75/4x + 3/4ky = -15x + 9y (-75+60/4)x+3/4ky=9y 15/4x+3/4ky=9y ____________________________________ So we have 45x + ky = 9y and in the end ky=9y/45x Now get 9y/45x to ky .

OpenStudy (anonymous):

-4/3?

OpenStudy (anonymous):

So we have this system : -15x+9y=-9 -20x + 9y / 45x = -12 Now can you answer to your self ?!:) :D

OpenStudy (anonymous):

12 is wrong. That's what i found first.

OpenStudy (anonymous):

alright, i'll follow the other guy's step and i'll see. thank you